OpenStudy (anonymous):

Can you check this? Samantha swam upstream for some distance in one hour. She then swam downstream the same river for the same distance in only 12 minutes. If the river flows at 4 mph, how fast can Samantha swim in still water? Is it distance = rate * time? d = 4r * 60t d = 4r * 12t I don't know how to find the mph in still water. Also, aren't I supposed to add/subtract something? It's a multiple choice but I don't want to guess wrong.

hartnn (hartnn):

yes, Is it distance = rate * time? thats correct but when she swims with the overall rate will be SUM of her rate and water and when she swims AGAINST water, the overall rate will be the difference makes sense ?

hartnn (hartnn):

swims with the water***

OpenStudy (anonymous):

So when she is going against it, would I change the equation to: d = -4r * 12t

OpenStudy (anonymous):

Wait that doesn't sound right. Do I subtract from something else in the equation?

hartnn (hartnn):

let take some variables her rate in still water = r rate of flow of water = w so overall rate when she goes with water will be r+w ok?

hartnn (hartnn):

and overall rate when she goes AGAINST water will be r-w got these ?

OpenStudy (anonymous):

Oh okay, I understand what you are talking about.

hartnn (hartnn):

so now again use the equation., distance = rate * time and tell me what u get ?

OpenStudy (anonymous):

Would I use these equations? d = 4 - w * 60t d = 4 + w * 12t

hartnn (hartnn):

where does 't' come from ?? t itself =60 in 1st case and 12 in other.... also preferable is to keep 't' in hours like t=1 in 1st case and t = 12/60 in 2nd

hartnn (hartnn):

and w = 4

hartnn (hartnn):

we need to find r

OpenStudy (anonymous):

the t was for time but if i take out the t, all i will have is: d = r - 4 d = r + 4

hartnn (hartnn):

see this: distance = rate * time for upstream, rate = r-w = r-4 time = 1 (in hours) so, d = (r-w)*1 got this ? can you make an equation for downstream ?

OpenStudy (anonymous):

d = (r-w)*1 would be the upstream and d = (r+w)*12/60 would be the downstream?

hartnn (hartnn):

correct! w=4 so can u get the value of r now ?

OpenStudy (anonymous):

would she be going at be 6mph?

OpenStudy (anonymous):

i mean i got 6 for the r

hartnn (hartnn):

let me check 60r-240 = 12r+48 48r = 288 yes! r= 6mph :D

OpenStudy (anonymous):

awesome, i submitted and it was right! thank you so much for your help!

hartnn (hartnn):

welcome ^_^